# Unsourced Random Access: A Comprehensive Survey

### v2

@@ IV-B URA Solutions for GMAC
As shown in the previous subsection, the conventional RA techniques are energy-inefficient for the URA setup. As a result, there have been substantial efforts to develop low-complexity and energy-efficient schemes that can be used in practical systems. The main approaches for this purpose in the literature can be categorized as slotting the transmission frame, coded compressed sensing (CCS), random spreading, sparse spreading, and on-off division MA (ODMA). In the following, we review the existing practical coding schemes for URA over GMAC based on this classification.
@@ IV-B1 Slotting-based schemes
One of the design strategies is slotting the transmission frame, where each active user randomly picks one or a few slots to transmit its data depending on the exact setting. In this direction, Ordentlich and Polyanskiy propose the first practical scheme for URA in [53], where a concatenated coding scheme is considered combined with the $T$-fold ALOHA protocol. $T$-fold ALOHA is an approach similar to slotted ALOHA in the sense that the users choose a random sub-block (slot) to transmit their packets, but in the slotted ALOHA, packets are lost in the case of a collision. However, in $T$-fold ALOHA, $T$ or fewer users can be simultaneously decoded in each slot. In [53], each user randomly chooses one slot to transmit its message. This concatenated coding scheme consists of an inner binary linear code which is used to decode the modulo-2 sum of the codewords transmitted within a slot and an outer code to recover the individual messages. Many off-the-shelf codes can be used for the inner code. The authors construct the outer code from the columns of a $T$-error correcting Bose–Chaudhuri–Hocquenghem (BCH) code. This scheme suffers in terms of energy efficiency as the performance gap with the random coding bound is about 20 dB when the number of active users is 300. Its performance is significantly improved in [54], where the authors employ a concatenated coding scheme with polar and BCH codes as the inner and outer codes, respectively, and propose a decoding approach exchanging the information between inner and outer decoding.
@@
The authors of [55] consider a slotted structure where the users can repeat their codewords across sub-blocks, which is referred to as T-fold IRSA. They propose to split the message into two parts and employ a coding scheme based on a combination of compressed sensing (CS) [56] and LDPC coding at the slot level. Namely, the users pick a preamble determined by the first part of the message from a common codebook, and the second part is encoded by an LDPC code and interleaved before the transmission. The preamble carries information about the interleaver and the utilized slots for a specific user. At the receiver side, the preamble part is recovered by the non-negative least squares (NNLS) algorithm and the coded part is decoded by joint belief propagation (BP). The replicas of the recovered codewords are peeled from the other slots via SIC. The general transmission structure in T-fold IRSA based URA solutions is depicted in Fig. 5.
$T$-fold IRSA is also utilized in [57] by replacing the LDPC codes with the polar codes and utilizing joint successive cancellation decoding, which provides a significant performance gain over [55]. In [58], the authors employ random signatures to convert the equal-gain channel into a more favorable i.i.d. fading scenario along with trellis-based codes in a $T$-fold ALOHA structure. Their proposed approach provides a low-complexity solution with moderate performance. Moreover, the authors of [59] employ finite-field coding in a slotted structure. More specifically, the users encode their data using a non-binary LDPC code and transmit it after repetition, zero-padding, and interleaving. At the receiver, intra-slot decoding is performed by a joint BP algorithm over a non-binary factor graph. Their proposed approach improves the performance of the slotted schemes in high user loads, compared to its previous counterparts.
@@ IV-B2 CCS-based schemes
Another main approach in URA over GMAC is called coded compressed sensing (CCS) [60, 61, 62, 63, 64, 65, 66], where the messages are divided into sub-blocks and each sub-block is first encoded by an outer code by appending redundant parity bits before being mapped to a CS codebook for the inner encoding. The encoded sub-blocks are then transmitted in different slots. At the receiver side, each sub-block is recovered via compressed sensing algorithms, and the recovered segments are stitched together by an outer tree code. The general encoding-decoding structure in coded compressed sensing is illustrated in Fig. 6. More specifically, the encoding process is performed in two phases: 1) In the tree encoder, $B$ bits of information are divided into $S$ parts, and all of these segments, except the initial one, are appended by additional parity-check bits; 2) in the CS encoder, each of $S$ sub-message is mapped to a CS codebook to select a sequence for transmission. Hence, the received signal is comprised of $S$ slots, with each slot being a superposition of the sequences from $K_{a}$ users. The decoder also functions in two stages: 1) In the initial stage, $K_{a}$ sequences associated with the $K_{a}$ active users are identified within each slot. 2) in the second stage, called outer decoder or tree decoder, with the assistance of the unique parity bits, the sub-messages corresponding to the detected sequences are combined to create a message of length $B$.
@@
Several variations of the CCS approach are examined in the URA literature. Specifically, in [60], the sub-blocks are first encoded by a binary linear code and then mapped to the columns of a common sensing matrix for inner encoding. In [61], sparse regression codes (SPARCs) are exploited as inner codes with approximate message passing (AMP) decoding at the receiver. The authors of [62] improve the performance of the proposed scheme in [61] by introducing a dynamic interaction between the inner AMP and the outer tree decoder. SPARCs are also employed in [63], where the sub-blocks are CRC-encoded and connected via block Markov superposition transmission. At the receiver, a hybrid decoder combining SIC and AMP is employed for inner decoding, with the addition of tree decoding. Furthermore, the authors of [64] adapt constant-weight codes as the outer code with the use of Gabor frames to design the inner CS coding matrix, and a concatenation of Raptor codes and tree codes is employed as the outer code in [65]. Finally, in a recent work [66], the authors propose a CCS-based scheme eliminating the tree code by using the encoded sub-message in the first slot as a user identifier. This reduces the parity bits and the resulting rate loss, offering the best performance among the CCS-based schemes on GMAC and outperforming the remaining solutions for $K_{a}\geq 225$.
@@ IV-B3 Random spreading-based schemes
In another design strategy for URA over GMAC, there is no slotting and the active users utilize the whole transmission frame. In some works following this strategy, the idea is to spread the channel-coded bits through the transmission time frame using random signatures. The random spreading idea was first introduced in [67], where the message is divided into two parts: the first part determines the signature sequence while the second part is encoded by a polar code. At the receiver, first, the selected signature sequences are detected by a correlation-based energy detector. Then, an iterative decoding algorithm consisting of MMSE filtering followed by single-user polar decoding is employed to recover the channel-coded bits. Moreover, SIC is applied at the end of each iteration. The general decoding process of random spreading-based solutions is illustrated in Fig. 7.
In [68], the authors modify the scheme in [67] by replacing the polar codes with polarization-adjusted convolutional (PAC) codes, with a slight performance improvement and reduced implementation complexity. Moreover, [44] improves the performance of the scheme in [67] by letting the users to randomly select a power level from a pre-defined set. This approach, called power diversity, creates an artificial fading-like scenario, where the users with higher power levels are decoded first with a high probability, and the ones with lower power levels are decoded next after SIC. This results in a performance improvement over [67] for $K_{a}\geq 150$. The authors of [69] replace the polar codes in [67] with LDPC codes to exploit the soft information produced by the belief propagation decoder to perform soft-input-soft-output (SISO) MMSE filtering. This approach improves the decoding performance since the information about the partially decoded codewords can be utilized, and it provides the best performance among the schemes based on random spreading. Furthermore, enhanced spread spectrum ALOHA (E-SSA) [70], an asynchronous RA protocol combining ALOHA with direct sequence spread spectrum is adapted for URA over GMAC in [71]. The authors employ a wrap-around version of E-SSA in order to adapt it for a synchronous URA setup.
@@ IV-B4 IDMA-based schemes
In random spreading based solutions, each active user occupies the entire time frame by random spreading of channel-coded bits for data transmission. In some other works, the active users distribute their codeword bits onto the transmission time frame in different ways. For instance, in the sparse interleave-division MA (IDMA) approach [72], the message bits and the time frame are divided into parts. The users transmit a preamble sequence dictated by the first part of their message in the preamble part of the frame. The rest of the bits are encoded by an LDPC code, zero-padded, and permuted by an interleaver determined by the preamble to be transmitted along the rest of the channel frame. This encoding process is illustrated in Fig. 8. At the receiver, after decoding the preamble part via the least absolute shrinkage and selection operator (LASSO) technique, a single joint Tanner graph is utilized for message recovery, exploiting the information of the location of transmitted symbols from the identified preamble.
@@
A similar approach to the sparse IDMA in [72] is adopted in [73] employing polar codes, however, different code lengths are used among the users and TIN is utilized as the decoding algorithm. Moreover, in [74], the message bits are first encoded by a high-rate error correction code. The encoded bits are divided into two parts, with the first part being mapped to a column of a common sensing matrix, which determines the interleaver and signature to be utilized in the second part after its repetition. The receiver utilizes AMP to recover the first part followed by a multiuser decoding algorithm for the second part. Furthermore, the authors of [75] propose to encode the data as the Kronecker product of two-component codewords, and employ an iterative decoder based on bilinear generalized AMP to decompose the Kronecker product and a soft-in-soft-out decoder for the individual components.
Table III: Main URA solutions for GMAC
Paper Approach Contribution and Techniques
@@
Polyanskiy, 2017 [11] - information-theoretic formulation of URA, and Random coding achievability and converse bounds
@@
Ordentlich et al., 2017 [53] slotting concatenated coding combined with $T$-fold ALOHA
Vem et al., 2019 [55] slotting LDPC coding, preamble transmission, and SIC
Marshakov et al., 2019 [57] slotting preamble transmission, polar coding, and SIC
@@
Amalladinne et al., 2020 [60] CCS inner CS and outer tree coding
Amalladinne et al., 2022 [62] CCS interaction between the inner AMP and the outer BP
Nassaji et al., 2024 [66] CCS eliminating the outer tree code
Pradhan et al., 2022 [72] sparse spreading preamble transmission, LDPC coding and SIC
Han et al., 2021 [75] sparse spreading convolutional coding and SIC
Pradhan et al., 2020 [67] random spreading polar coding and SIC
Ahmadi et al., 2020 [44] random spreading power diversity, polar coding, and SIC
Pradhan et al., 2021 [69] random spreading LDPC coding, soft cancellation, and SIC
Yan et al., 2023 [76] ODMA repeat-accumulate coding and soft cancellation
Ozates et al., 2024 [78] ODMA polar coding and SIC
Yan et al., 2024 [79] ODMA non-binary LDPC coding and soft cancellation
@@ IV-B5 ODMA-based schemes
Another way to distribute the user codeword bits onto the transmission time frame is proposed in [76], where the authors employ on-off division MA (ODMA) [77] in the context of URA. In this approach, the active users exploit a small disjoint portion of the transmission time frame to transmit their data encoded by a repeat-accumulate code according to a randomly selected transmission pattern, with the remaining time instances being idle. At the receiver, the BS first recovers the transmission patterns and then employs multiuser detection and decoding over a sparse factor graph. The authors of [78] investigate the ODMA technique in URA further, namely, they employ polar codes to improve the performance in the low active user load regimes combined with a simpler pattern detection algorithm. Another high-performing solution is proposed in [79], which improves the scheme in [76] by employing a non-binary LDPC code.
@@
IV-C Summary and Performance Comparison
A summary of main URA solutions for GMAC can be found in Table III, accompanied by a comparison of their performance in terms of energy efficiency in Fig. 9 for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature. The results in Fig. 9 demonstrate that random spreading-based schemes in [67, 44, 69] offer good performance for a lower number of active users; however, their performance degrades with the increasing number of active users since each active user occupies the entire transmission frame. On the other hand, ODMA-based approaches in [76, 78, 79] outperform the others except the solution proposed in [66] in the high multiuser interference regime due to its sparsity, while it performs well in the low multiuser regime when combined with polar coding. Moreover, CCS-based approaches can exhibit very good performance for a high number of active users when the rate loss due to the redundant parity bits is alleviated, as shown in [66]. Compared to the achievability bound, the proposed schemes can perform similar or slightly better for $K_{a}\leq 100$, and the gap is a maximum of 0.8 dB for $K_{a}\leq 300$.
IV-D Computational Complexity
In terms of complexity, slotting is the most beneficial approach among the existing solutions for GMAC URA, as the complexity of the operations depending on the user count super-linearly increases by the number of active users in the slot rather than $K_{a}$. In addition, slotting leads to smaller preamble/pilot length and channel code length $n_{c}$, decreasing the complexity of the operations depending on these parameters. CCS also has a low complexity since slot-wise compressed sensing and a low-complexity tree code are employed in conjunction. Namely, its complexity order is linear or quadratic in $K_{a}$ depending on the regime [60]. ODMA is another low-complexity approach, as only a small fraction of the time frame is occupied by each user, which simplifies the iterative multiuser decoding process [76]. More specifically, its complexity order linearly depends on $n_{c}$ and $K_{a}$. On the other hand, random spreading-based schemes where each user occupies the entire time frame have the highest complexity, as the complexity order is cubic in $K_{a}$ due to the MMSE filtering [67]. On another note, the number of computations needed for some decoding operations depends on the transmission time frame length rather than the slot or code length as in slotting or ODMA, which increases the complexity.

### v3

@@ IV-B URA Solutions for GMAC
As shown in the previous subsection, the conventional RA techniques are energy-inefficient for the URA framework. As a result, there have been substantial efforts to develop low-complexity and energy-efficient schemes that can be used in practical systems. The main approaches for this purpose in the literature can be categorized as slotting the transmission frame, coded compressed sensing (CCS), random spreading, interleave-division MA (IDMA), and on-off division MA (ODMA). In the following, we review these schemes.
@@ IV-B1 Slotting-based schemes
One of the design strategies is slotting the transmission frame, where each active user randomly picks one or a few slots to transmit its data, depending on the exact setting. In this direction, Ordentlich and Polyanskiy propose the first practical scheme for URA in [55], where a concatenated coding scheme combined with the $T$-fold ALOHA protocol is considered. $T$-fold ALOHA is an approach similar to slotted ALOHA in the sense that the users choose a random sub-block (slot) to transmit their packets, but in slotted ALOHA, packets are lost in the case of a collision. However, in $T$-fold ALOHA, $T$ or fewer users can be simultaneously decoded in each slot. In [55], each user randomly chooses one slot to transmit its message. This concatenated coding scheme consists of an inner binary linear code, which is used to decode the modulo-2 sum of the codewords transmitted within a slot, and an outer code to recover the individual messages. Many off-the-shelf codes can be used for the inner code. The authors construct the outer code from the columns of a $T$-error correcting Bose–Chaudhuri–Hocquenghem (BCH) code. This scheme suffers in terms of energy efficiency as the performance gap with the random coding bound is about 20 dB when the number of active users is 300. Its performance is significantly improved in [56] (by up to 18 dB), where the authors employ a concatenated coding scheme with polar and BCH codes as inner and outer codes, respectively, and propose a decoding approach that exchanges information between inner and outer decoders. Specifically, all the candidate codewords from the inner successive cancellation list decoding of the polar code, rather than just one, are passed to the outer BCH decoder to facilitate its error checking. In addition, indicators for the decoding status of the outer BCH decoder are used to enhance the inner polar decoding process.
@@
The authors in [57] consider a slotted structure where the users can repeat their codewords across sub-blocks, which is referred to as $T$-fold IRSA. They propose splitting the message into two parts and employing a coding scheme based on a combination of compressed sensing (CS) [58] and LDPC coding at the slot level. Namely, the users select a preamble determined by the first part of the message from a common codebook, and the second part is encoded using an LDPC code and interleaved before transmission. The preamble carries information about the interleaver and the utilized slots for a specific user. At the receiver side, the preamble part is recovered by a non-negative least squares (NNLS) algorithm, and the coded part is decoded by joint belief propagation (BP). The replicas of the recovered codewords are peeled from the other slots via SIC. This scheme improves the performance of [55] by up to 13 dB, and the difference with the random coding bound is less than 8 dB. The general transmission structure in $T$-fold IRSA based URA solutions is depicted in Fig. 6, where $N$ is the length of the transmission frame, $N_{s}$ is the slot length, $V$ is the number of slots, $w_{i}$ is the message of the $i$-th user, and $x(w_{i})$ denotes the transmitted signal of the $i$-th user.
$T$-fold IRSA is also utilized in [59] by replacing the LDPC codes with polar codes and utilizing joint successive cancellation decoding [60], which provides a significant performance gain over [57] due to the superiority of the polar codes to the LDPC codes in the short blocklength regime. In [61], the authors employ random signatures to convert the equal-gain channel into a more favorable i.i.d. fading scenario along with trellis-based codes in a $T$-fold ALOHA structure. Their proposed approach provides a low-complexity solution with moderate performance. Moreover, the authors in [62] employ finite-field coding in a slotted structure. More specifically, the users encode their data using a non-binary LDPC code and transmit it after repetition, zero-padding, and interleaving. At the receiver, intra-slot decoding is performed using a joint belief propagation (BP) algorithm over a non-binary factor graph. Their proposed approach improves the performance of the previously proposed slotted URA schemes for high active user loads.
@@ IV-B2 CCS-based schemes
Another main approach in URA over GMAC is called coded compressed sensing (CCS) [63, 64, 65, 66, 67, 68, 69], where the messages are divided into sub-blocks and each sub-block is first encoded by an outer code by appending parity bits before being mapped to a CS codebook for inner encoding. The encoded sub-blocks are then transmitted in different slots. At the receiver side, each sub-block is recovered via compressed sensing algorithms, and the recovered segments are stitched together by an outer code, where a tree code is typically used. The general encoding-decoding structure in coded compressed sensing is illustrated in Fig. 7, where $G$ is a random binary matrix performing tree encoding, $v_{i}$ is the output signal of the tree encoder for the $i$-th user, and $\hat{W}$ is the set of decoded messages. More specifically, the encoding process is performed in two phases: 1) In the tree encoder, $B$ bits of information are divided into $S$ parts, and all of these segments, except the initial one, are appended additional parity-check bits; 2) in the CS encoder, each $S$ sub-message is mapped to a CS codebook to select a sequence for transmission. Hence, the received signal is comprised of $S$ slots, with each slot being a superposition of the sequences from $K_{a}$ users. Note that no parity-check bits are appended to the initial segment of the tree encoder since this segment is used as a root node in the tree decoding process [63]. The decoder also functions in two stages: 1) In the initial stage, $K_{a}$ sequences associated with the $K_{a}$ active users are identified within each slot. 2) In the second stage, called outer decoder or tree decoder, with the assistance of the unique parity bits, the sub-messages corresponding to the detected sequences are combined to create a message of length $B$.
@@
Several variations of the CCS approach are examined in the URA literature. Specifically, in [63], the sub-blocks are first encoded by an outer tree code and then mapped to the columns of a common sensing matrix for inner encoding. In [64], sparse regression codes (SPARCs), a class of channel codes that can achieve channel capacity of a point-to-point AWGN channel under approximate message passing (AMP) decoding, are exploited as inner codes to improve the performance of the scheme in [63], while the outer code is again a tree code. The authors in [65] further improve the performance of the approach in [64] by introducing a dynamic interaction between the inner AMP and the outer tree decoder. SPARCs are also employed in [66], where the sub-blocks are CRC-encoded and connected via block Markov superposition transmission. At the receiver, a hybrid decoder combining SIC and AMP is used for inner decoding, with the addition of tree decoding. Furthermore, the authors in [67] employ constant-weight codes as outer codes with the use of Gabor frames to design the inner CS coding matrix, and a concatenation of Raptor codes and tree codes is used as the outer code in [68]. Finally, in a recent work [69], the authors propose a CCS-based scheme eliminating the tree code by using the encoded sub-message in the first slot as a user identifier. This reduces the parity bits and the resulting rate loss, offering the best performance among the CCS-based schemes on GMAC and outperforming the other solutions for $K_{a}\geq 225$.
@@ IV-B3 Random spreading-based schemes
In another design strategy for URA over GMAC, there is no slotting, and the active users utilize the entire transmission frame. In some works that follow this strategy, the idea is to spread the channel-coded bits throughout the transmission frame using random signatures. The random spreading idea was first introduced in [70], where the message is divided into two parts: the first part determines the signature sequence, while the second part is encoded by a polar code. At the receiver, first, the selected signature sequences are detected by a correlation-based energy detector. Then, an iterative decoding algorithm consisting of MMSE filtering followed by single-user polar decoding is employed to recover the channel-coded bits. Moreover, SIC is applied at the end of each iteration. The general decoding process of random spreading-based solutions is illustrated in Fig. 8, where $\mathbf{Y}$ is the received signal, $\mathbf{Y}_{t}$ is the residual of the received signal at the $t$-th iteration, and $D^{t}$, $\beta^{t}_{i}$, and $\hat{v}^{t}_{i}$ are the set of detected signature indices, LLR vector, and the message estimate of the $i$-th user at the $t$-th iteration, respectively.
In [71], the authors modify the scheme in [70] by replacing the polar codes with polarization-adjusted convolutional (PAC) codes, with a slight performance improvement and reduced implementation complexity. The performance of the scheme in [70] is improved in [44] by allowing users to randomly select a power level from a predefined set. This approach, called power diversity, creates an artificial fading-like scenario, where the users with higher power levels are decoded first with a high probability, and those with lower power levels are decoded next after SIC. This results in a performance improvement over [70] for $K_{a}\geq 150$. In [72], the authors propose to encode the data as the Kronecker product of two-component codewords, and employ an iterative decoder based on bilinear generalized AMP to decompose the Kronecker product and a soft-in-soft-out decoder for the individual components. The authors in [73] replace the polar codes in [70] with LDPC codes to exploit the soft information produced by the belief propagation decoder to perform soft-input-soft-output (SISO) MMSE filtering. This approach enhances decoding performance by leveraging information from partially decoded codewords, resulting in the best performance among schemes based on random spreading. Moreover, enhanced spread spectrum ALOHA (E-SSA) [74], an asynchronous RA protocol that combines ALOHA with direct sequence spread spectrum, is adapted for URA over GMAC in [75], where a wrap-around version of E-SSA is employed to adapt it for a synchronous URA setup. In a recent work, random spreading with higher order pulse amplitude modulation (PAM) such as 8-PAM or 16-PAM is studied in [76]. The authors propose an AMP-based solution combined with expectation maximization and multiple measurement vectors, demonstrating that their proposed solution can support up to 1100 active users. This is the highest number of supported users reported in the literature for URA over GMAC for the practical PUPE value of $\epsilon=0.05$.
@@ IV-B4 IDMA-based schemes
In random spreading-based solutions, each active user occupies the entire frame by random spreading of channel-coded bits for data transmission. In some other works, active users distribute their codeword bits across the transmission frame in various ways. For instance, in the sparse interleave-division MA (IDMA) approach [77], the message bits and the frame are divided into different parts. The users transmit a preamble sequence dictated by the first part of their message in the preamble part of the frame. The remaining bits are encoded using an LDPC code, zero-padded, and permuted by an interleaver determined by the preamble to be transmitted over the rest of the channel frame. This encoding process is illustrated in Fig. 9. At the receiver, after decoding the preamble part using the least absolute shrinkage and selection operator (LASSO), a single joint Tanner graph is employed for message recovery, leveraging the information on the location of transmitted symbols obtained from the identified preamble. In a recent sparse IDMA-based scheme proposed in [78], the authors employ AMP for the recovery of the preamble part and convolutional coding in the data part, which improves the performance of [77] in high active user loads.
@@
A similar approach to the sparse IDMA in [77] is adopted in [79], employing polar codes; however, different code lengths are used among the users, and TIN is utilized as the decoding algorithm. Moreover, in [80], the message bits are first encoded by a high-rate error correction code. The encoded bits are divided into two parts, with the first part being mapped to a column of a common sensing matrix, which determines the interleaver and signature to be utilized in the second part after its repetition. The receiver utilizes AMP to recover the first part, followed by a multiuser decoding algorithm for the second part.
While the works mentioned above consider the synchronous transmission, a fully asynchronous (i.e., with an unbounded time horizon) URA setup is studied in [81], where each user transmits a preamble (from a pool of available ones) for timing acquisition followed by its payload that is interleaved and multiplied by a signature similar to [80]. The receiver employs a sliding window decoding approach where correlation-based arrival time detection, multiuser decoding, and SIC operations are iteratively performed.
Table III: A summary of URA solutions for GMAC
Paper Approach Description Channel Code Synchron. vs. Asynch.
@@
Polyanskiy, 2017 [11] - Information-theoretic formulation of URA, and Random coding achievability and converse bounds. - -
@@
Ordentlich et al., 2017 [55] slotting $T$-fold ALOHA-based concatenated coding scheme. BCH Synchron.
Vem et al., 2019 [57] slotting $T$-fold IRSA-based scheme employing preamble transmission, channel coding, joint BP decoding, and SIC. LDPC Synchron.
Marshakov et al., 2019 [59] slotting $T$-fold IRSA-based scheme employing preamble transmission, channel coding, joint successive cancellation decoding, and SIC. Polar Synchron.
@@
Amalladinne et al., 2020 [63] CCS Concatenated coding scheme employing inner CS and outer tree coding. - Synchron.
Amalladinne et al., 2022 [65] CCS Concatenated coding scheme employing inner SPARC and outer tree codes. Interaction between inner AMP and outer BP decoding is proposed. - Synchron.
Nassaji et al., 2024 [69] CCS Concatenated coding scheme employing SPARCs and CRC encoding. Eliminating the outer tree code is proposed to improve the performance. - Synchron.
Pradhan et al., 2020 [70] random spreading Spreading-based scheme combining correlation-based signature detection, MMSE symbol estimation, single-user decoding, and SIC. Polar Synchron.
Ahmadi et al., 2020 [44] random spreading Spreading-based scheme employing similar operations with [70] and improve its performance with power diversity. Polar Synchron.
Han et al., 2021 [72] random spreading Spreading-based scheme combining Bayesian matrix factorization, BCJR decoding, and SIC at the receiver side. Convolutional Synchron.
Pradhan et al., 2021 [73] random spreading Spreading-based scheme combining energy detection, soft MMSE estimation, soft cancellation, single-user decoding, and SIC. LDPC Synchron.
Pradhan et al., 2022 [77] IDMA Spreading-based scheme employing preamble transmission, joint Tanner graph decoding, and soft cancellation. LDPC Synchron.
Karami et al., 2024 [81] IDMA Spreading-based scheme employing correlation-based timing detection, multiuser decoding and SIC. - Asynch.
Yan et al., 2023 [82] ODMA ODMA-based scheme employing blind pattern detection, joint graph decoding, and soft cancellation. RA Synchron.
Ozates et al., 2024 [84] ODMA ODMA-based scheme employing energy test for pattern detection, single-user decoding, and SIC. Polar Synchron.
Yan et al., 2024 [85] ODMA ODMA-based scheme employing blind pattern detection, joint graph decoding, collision resolution, and soft cancellation. Non-binary LDPC Synchron.
Ozates et al., 2025 [87] ODMA ODMA-based scheme employing nested sliding windows, timing detection by ODMA patterns, single-user decoding and SIC. Polar Asynch.
@@ IV-B5 ODMA-based schemes
Another way to distribute the user codeword bits onto the transmission frame is proposed in [82], where the authors employ on-off division MA (ODMA) [83] in the context of URA. In this approach, the active users exploit a small disjoint portion of the transmission frame to transmit their data encoded by a repeat-accumulate code according to a randomly selected transmission pattern, with the remaining time instances being idle. At the receiver, the BS first recovers the transmission patterns and then employs multiuser detection and decoding over a sparse factor graph. The authors in [84] investigate the ODMA technique in URA further, namely, they employ polar codes to improve the performance in the low active user load regimes combined with a simpler pattern detection algorithm. Another high-performing solution is proposed in [85], which improves the scheme in [82] by employing a non-binary LDPC code. Moreover, an analytical power division method that is formulated based on TIN-SIC is proposed in [86], which achieves the same performance as that of [85] using 5G-NR LDPC codes instead of non-binary LDPC coding. Furthermore, the authors in [87] examine the fully asynchronous setup introduced in [81], and propose a preamble-free ODMA-based solution. The receiver utilizes a sliding-window decoding approach with two nested windows. In the inner sliding window, an iterative decoder combining arrival time detection using the ODMA patterns, single-user polar decoding, and SIC is employed, while the outer window enhances the interference cancellation.
@@
IV-C Summary and Comparison of GMAC Solutions
A summary of the URA solutions for GMAC is presented in Table III, with comparisons among them in different aspects.
IV-C1 Performance Comparison
The performance of the GMAC URA schemes is compared in terms of energy efficiency, the standard performance criterion in URA, in Fig. 10 for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature. Note that due to the vast number of papers, we have included only the state-of-the-art papers, at least in some regime, and those introducing a new approach when published. The results in Fig. 10 demonstrate that random spreading-based schemes in [70, 44, 73] offer good performance for a low number of active users; however, their performance degrades with increasing the number of active users since each active user occupies the entire transmission frame. On the other hand, ODMA-based approaches in [82, 84, 85] outperform the others, except for the solution proposed in [69], in the high multiuser interference regime due to their sparsity. Meanwhile, they perform well in the low multiuser regime when combined with polar coding. Moreover, the CCS-based approaches can exhibit very good performance for a high number of active users when the rate loss due to the parity bits is alleviated, as shown in [69]. Compared to the achievability bound, the proposed schemes can perform similarly or slightly better for $K_{a}\leq 100$, and the gap is at most about 0.8 dB for $K_{a}\leq 300$.
In summary, the performance trends over GMAC reveal a clear trade-off: While random spreading-based schemes are attractive at low user loads, their scalability is limited due to full-frame occupancy and high complexity. Conversely, ODMA-based approaches and recent CCS variants achieve robust performance in high-user regimes, benefiting from sparsity and efficient decoding. These results underscore that practical URA deployment must carefully balance energy efficiency, complexity, and user density, with CCS and ODMA schemes emerging as the most scalable options.
IV-C2 Computational Complexity Comparison
In terms of complexity, slotting is the most beneficial approach among the existing solutions for URA over GMAC, as the complexity of the operations, depending on the user count, increases super-linearly by the number of active users in the slot rather than $K_{a}$. In addition, slotting leads to smaller preamble/pilot lengths and a shorter channel code length $n_{c}$, thereby decreasing the complexity of the operations. Coded compressed sensing schemes also offer low complexity since slot-wise compressed sensing and low-complexity tree codes are employed in conjunction with each other. Specifically, its complexity order is linear or quadratic in $K_{a}$, depending on the regime [63]. ODMA is another low-complexity approach, as only a small fraction of the frame is occupied by each user, which simplifies the iterative multiuser decoding process [82]. More specifically, its complexity order depends linearly on $n_{c}$ and $K_{a}$. On the other hand, random spreading-based schemes where each user occupies the entire frame have the highest complexity, as the complexity order is cubic in $K_{a}$ due to the MMSE filtering [70].